# Digital Twin-Empowered Task Assignment in Aerial MEC Network: A Resource Coalition Cooperation Approach with Generative Model

### v2

@@ Abstract
To meet the demands for ubiquitous communication and temporary edge computing in 6G networks, aerial mobile edge computing (MEC) networks have been envisioned as a new paradigm. However, dynamic user requests pose challenges for task assignment strategies. Most of the existing research assumes that the strategy is deployed on ground-based stations or UAVs, which will be ineffective in an environment lacking infrastructure and continuous energy supply. Moreover, the resource mutual exclusion problem of dynamic task assignment has not been effectively solved. Toward this end, we introduce the digital twin (DT) into the aerial MEC network to study the resource coalition cooperation approach with the generative model (GM), which provides a preliminary coalition structure for the coalition game. Specifically, we propose a novel network framework that is composed of an application plane, a physical plane, and a virtual plane. After that, the task assignment problem is simplified to convex optimization programming with linear constraints. And then, we also propose a resource coalition cooperation approach that is based on a transferable utility (TU) coalition game to obtain an approximate optimal solution. Numerical results confirm the effectiveness of our proposed approach in terms of energy consumption and utilization of resources.
@@ I Introduction
In order to better introduce the motivation of this paper, we summarize existing task execution methods into the following four categories. An example in Fig. 1 is given below to illustrate the challenges that existing multi-UAV systems deal with task offloading. In Fig. 1 a), tasks are offloaded to UAVs randomly or selectively without considering the UAVs’ actual computing, storage, and communication resources. This often results in poor task offloading service quality and inefficient resource utilization due to resource shortages or surpluses. Fig. 1 b) illustrates the use of the equilibrium method among multi-UAV providing one-to-one offloading services. Each UAV provides task offloading services independently, requiring careful evaluation of the inter-UAV relationships. However, this method leads to unjust resource distribution and, subsequently, lower task offloading service quality. Fig. 1 c) shows a scenario where all UAVs cooperate with each other to form a multi-UAV system to provide all-to-one task offloading services. The approach enhances the quality of task offloading services, it can also lead to underutilization of UAV resources. In contrast, Fig. 1 d) differs from Fig. 1 c) by simultaneously considering task offloading service requests and resource cooperation. Multiple UAVs unite to form a team with the aim of providing many-to-one task offloading services while maximizing overall utility. This method excels in optimizing both resource utilization and task offloading service quality. We clearly see that the method described in Fig. 1 d) is more in line with actual needs. However, the UAV-assisted MED also faces new challenges, i.e., the long line-of-sight link between the user and the UAV will cause a long transmission latency, and the computing resources at the UAV are not always adequate, which are not friendly to delay-sensitive tasks [9]. Although some excellent work has demonstrated that device-to-device task offloading can be used to ensure low latency and high data transmission rate [10, 11, 12], the mobility and service requirements of user and the unpredictability of UAV resources in the real world have not yet been fully considered.
@@
Inspired by the motivation and the emerging technique, this paper considers DT-empowered task assignment with resource coalition cooperation in an aerial MEC network consisting of multiple users, UAVs equipped with edge servers, and an airship equipped with a more powerful cloud server, where the users randomly generate tasks as they move. In this case, our goal is to reduce energy consumption by jointly optimizing the task assignment, energy consumption, communication bandwidth, and the resource utilization. For the sake of clarity, the main contributions of this work are listed as follows
@@
The rest of the section-wise paper outline is detailed as follows. Section II presents a brief synopsis of recent literature in this area. Section III introduces the system model. The question statement of the TU coalition is presented in Section IV. A novel TU coalition game algorithm is given in Section V. The simulation results are discussed in Section VI and conclude this paper in Section VII.
@@ II-C Application of GM in Edge Intelligence Network
To our knowledge, this is the first time we have researched DT-empowered task assignment by coalition game with a GM in an aerial MEC network. Specifically, the system model can monitor the status of the network in real-time and use the method of resource coalition cooperation to optimize task assignment and execution efficiency.
III System Model
This system is designed to use high-altitude airships with high computing power, large caches, and strong communications to assist multi-UAV in completing resource optimization to adapt to MED’s task assignment requests. We build DT models of UAVs in an airship to optimize the resource cooperation of aerial MEC network. With the help of DT, the system model designed in this paper is shown in Fig. 2, which is composed of a application plane, a physical plane, and a virtual plane. The model parameters are shown in Table I. Next, we describe the system model as follows.
@@ III System Model
III-A DT Model
In this paper, the communication between MED and UAV (M2U) is designed to obtain the MED task parameters to build a DT model. In addition, the communication between UAV and airship (U2A) is designed to transmit the DT modeling parameters of MED and UAV. The above communication process all adopts the communication method of time division multiple access. To simplify the analysis, this paper defines the system model as a synchronous periodic time slot system, and the duration of each time slot $t$ is $\tau$. Let the MED set be $M=\left\{{1,2,\cdots,i,\cdots,m}\right\}$. The $i$th MED is represented as $MED_{i}$. Then the DT of $i$th MED in time slot $t$ is
$DT^{t}_{MED_{i}}=\{l^{t}_{MED_{i}},s^{t}_{i},\vartheta^{t}_{i},p^{t}_{i,tr},\tau^{t}_{i}\}$ (1)
where $l^{t}_{MED_{i}}$ represents coordinate $(x^{t}_{i},y^{t}_{i},z^{t}_{i})$ of $MED_{i}$. $s^{t}_{i}$ is the amount of tasks that $MED_{i}$ requests to assign within the time slot $t$. $\vartheta^{t}_{i}$ represents the computing resource required to process the amount of tasks per unit, that is, computational complexity. $p^{t}_{i,tr}$ represents the transmission power of $MED_{i}$ for task assignment. $\tau^{t}_{i}$ represents the maximum time allowed to complete task of $MED_{i}$, which is $\tau^{t}_{i}\leq\tau$.
The set of UAV is denoted as $V=\left\{{1,2,\cdots,j,\cdots,n}\right\}$. Let $UAV_{i}$ be the $i$th UAV, DT of $UAV_{j}$ within time slot $t$ is given as follows
$DT^{t}_{UAV_{j}}=\{l^{t}_{UAV_{j}},b^{t}_{j},f^{t}_{j},c^{t}_{j},p^{t}_{j,h}\}$ (2)
@@ III-A DT Model
TABLE I: List of parameters
@@
III-B Communication Model
Assume that only one $MED_{i}$ interacts with the UAV coalition $N$ in each time slot $t$, and the coordinated multiple points transmission technology based on orthogonal frequency division multiple access is introduced in the UAV coalition. And allows UAV to reuse spectrum resources with a total bandwidth of $B^{t}_{max}$ to reduce communication interference. At the same time, other UAVs and MED outside the UAV coalition are in energy-saving mode. Further, the communication capacity between $MED_{i}$ and $UAV_{j}$ in the time slot $t$ is given as follows
$C^{t}_{ij}=b^{t}_{j}\log\left(1+\dfrac{|g^{t}_{ij}|p^{t}_{i,tr}}{\sigma^{2}}\right)$ (3)
@@ III-B Communication Model
III-C Energy Consumption Model
@@ IV The Question Statement of Task Assignment
In this section, we first introduce the task assignment problem in the aerial MEC network and discuss its unique challenges compared with conventional task offloading in MEC networks. We then formulate a new task assignment problem in the network, which concentrates on improving the energy consumption and utilization of resources.
IV-A Definition and Challenges of Task Assignment
The cooperation of resources from various computing devices is a fundamental problem in MEC network. The basic idea of task assignment is to offload the tasks to the appropriate computing devices, which is to improve the utilization of resources and the efficiency of task execution. However, in order to find the optimized task assignment strategy, we are mapping the multiple UAVs to the resource pool. Compared with the task offloading in MEC systems, the task assignment problem has its unique challenges. Consequently, we provide analysis toward the unique characteristics of task assignment in aerial MEC network.
Peer-to-peer pooling: All the devices equipped with various resources in the aerial MEC network are regarded as equal network nodes, which build a resource pool of MEC network. Specifically, the resources of all nodes can perform tasks collaboratively through cooperative games. In other words, these nodes are treated equally as the computing devices, and their resource capabilities will be uniformly scheduled.
Low-to-high scheduling: The nodes in the aerial MEC network are composed of end users, edge servers, and cloud servers. The tasks in our proposed network can be executed in low-to-high manner according to task request: 1) one-to-one executing; 2) one-to-many assigning. In this paper, we focus on scheduling resources on multiple nodes to complete user task requests.
One-to-many assigning: In existing task offloading strategies, the tasks are usually offloaded to a powerful MEC server or a remote cloud server for increasing computing efficiency (i.e., one-to-one executing). However, in the proposed aerial MEC network, the tasks can be assigned from one node to many nodes.
@@ IV-C TU Coalition Utility
$\begin{array}[]{ll}F1:&\underset{s^{t}_{j},b^{t}_{j}}{\text{max}}~{}~{}~{}U^{t}_{g_{k}}\\ \text{ s.t.: }&C_{1}:0\leq b^{t}_{j}\leq b^{t}_{j,max}\\ &C_{2}:\Sigma_{j\in g_{k}}b^{t}_{j}\leq B^{t}_{max}\\ &C_{3}:0\leq s^{t}_{j}\leq c^{t}_{j,max}\\ &C_{4}:\Sigma_{j\in g_{k}}s^{t}_{j}\leq s^{t}_{i}\\ &C_{5}:T^{t}_{gk}\leq\tau^{t}_{i}\end{array}$ (9)
@@ IV-D Coalition Participant Utility
$u_{j}^{t}=U^{t}_{g_{k}}\frac{s_{j}^{t}}{\sum_{j\in g_{k}}s_{j}^{t}}-\alpha E_{j,cp}^{t}-\beta E_{j,h}^{t}$ (10)
@@
$\begin{array}[]{ll}F2:&\underset{f^{t}_{j}}{\text{max}}~{}~{}~{}u^{t}_{j}\\ \text{ s.t.: }&\mathrm{C}_{6}:0\leq f_{j}^{t}\leq f^{t}_{j,max}\\ &\mathrm{C}_{7}:T_{j,cp}^{t}\leq\tau_{i}^{t}-T_{j,tr}^{t}\end{array}$ (11)
@@ IV-E Problem Conversion and Merging
$\mathrm{C}_{8}:s_{j}^{t}\leq\tau_{i,\max}^{t}/\left(\frac{1}{C_{ij}^{t}}+\frac{\vartheta_{j}^{t}}{f_{j,\text{ max }}^{t}}\right)$ (13)
@@
$\begin{array}[]{ll}F3:&\underset{s_{j}^{t},b_{i}^{t}}{\text{max}}~{}~{}~{}U^{t}_{g_{k}}\\ \text{ s.t.: }&\mathrm{C}_{1},\mathrm{C}_{2},\mathrm{C}_{3},\mathrm{C}_{4},\mathrm{C}_{8},\mathrm{C}_{9}\end{array}$ (15)

### v3

@@ Abstract
To meet the demands for ubiquitous communication and temporary edge computing in 6G networks, aerial mobile edge computing (MEC) networks have been envisioned as a new paradigm. However, dynamic user requests pose challenges for task assignment strategies. Most of the existing research assumes that the strategy is deployed on ground-based stations or UAVs, which will be ineffective in an environment lacking infrastructure and continuous energy supply. Moreover, the resource mutual exclusion problem of dynamic task assignment has not been effectively solved. Toward this end, we introduce the digital twin (DT) into the aerial MEC network to study the resource coalition cooperation approach with the generative model (GM), which provides a preliminary coalition structure for the coalition game. Specifically, we propose a novel network framework that is composed of an application plane, a physical plane, and a virtual plane. After that, the task assignment problem is simplified to convex optimization programming with linear constraints. And then, we also propose a resource coalition cooperation approach that is based on a transferable utility (TU) coalition game to obtain an approximate optimal solution. Extensive analysis and numerical results confirm the effectiveness of our proposed approach in terms of energy consumption and utilization of resources.
@@ I Introduction
TABLE I: Comparison between related works and our work
Aerial MEC network architecture Optimization objective
Ref. One-to-many offloading HAP-assistance UAVs cooperation U2U interaction Energy consumption Resource utilization
[9]
[10]
[11]
[12]
[13]
[14]
[15]
[16]
[17]
[18]
Our work
@@
In order to better introduce the motivation of this paper, we summarize existing task execution methods into the following four categories. An example in Fig. 1 is given below to illustrate the challenges that existing multi-UAV systems deal with task offloading. In Fig. 1 a), tasks are offloaded to UAVs randomly or selectively without considering the multi-UAV cooperation. This often results in poor task offloading service quality and inefficient resource utilization due to resource shortages or surpluses. Related to this task offloading schedule under the multi-UAV system, in [9], the authors consider a multi-UAV-assisted MEC system. Each user could generate a task that could be offloaded to a UAV. And the authors propose a joint task offloading, computation resource allocation, and UAV trajectory control approach to solve a multi-objective optimization problem. In [10], the authors design the task for each user, which can be divided into two subtasks. Each user can execute one of its subtasks locally and offload the other to a UAV or ground-based station. The authors propose a joint scheme for UAV deployment and resource allocation when performing task offloading. Fig. 1 b) illustrates the use of the equilibrium method among multi-UAV providing one-to-one offloading services. Each UAV provides task offloading services independently, requiring careful evaluation of the inter-UAV relationships. However, this method leads to unjust resource distribution and, subsequently, lower task offloading service quality. In [11], the authors propose a cooperative reconnaissance task and spectrum access scheme for coalition-based UAV networks and prove that the proposed scheme can converge to a stable coalition while maximizing network utility. In [12], the authors propose UAV-assisted emergency communication with nonorthogonal multiple access and present a joint offloading decision and resource allocation problem to minimize computation overhead. Fig. 1 c) shows a scenario where all UAVs cooperate with each other to form a multi-UAV system to provide all-to-one task offloading services. The approach enhances the quality of task offloading services, it can also lead to underutilization of UAV resources. The representative scenario of [19] is similar to this case, all the cellular users cooperate with each other for mobile data forwarding task. In [13], the authors design a multi-UAV-assisted MEC system with software-defined networking technology to provide task offloading services for edge user devices within the region. The aim is to optimize UAV flight trajectories and minimize overall system energy consumption. In [14], the authors propose two possible sources of the tasks in practical scenarios, i.e., ground users and UAVs, and investigate the delay-aware cooperative task offloading problem for the multi-UAV-enabled edgecloud computing system. In contrast, Fig. 1 d) differs from Fig. 1 c) by simultaneously considering task offloading service requests and resource cooperation. Multiple UAVs to form a team with the aim of providing many-to-one task offloading services while maximizing overall utility. This method excels in optimizing both resource utilization and task offloading service quality. We clearly see that the method described in Fig. 1 d) is more in line with actual needs. In [15], the authors investigate overlapping coalition formation in a task-driven cooperative UAV network. The resources of UAVs are allocated to the same task to form a coalition to execute the task cooperatively. Numerical results show that the proposed algorithm increases the average utility of tasks.In [16], the authors propose an edge computing network architecture for space-air-ground integration and utilize a binary integer linear programming model to resolve the offloading cost minimization problem. Furthermore, they prove the existence of Nash equilibrium through the potential game. However, the UAV-assisted MED also faces new challenges, i.e., the long line-of-sight link between the user and the UAV will cause a long transmission latency, and the computing resources at the UAV are not always adequate, which are not friendly to delay-sensitive tasks [20]. Although some excellent work has demonstrated that device-to-device task offloading can be used to ensure low latency and high data transmission rate [21, 22, 23], the service requirements of user and the unpredictability of UAV resources in the real world have not yet been fully considered. And it can also be seen from the above four examples that the task assignment in a aerial UAVs is typically a dynamic process. The real-time task assignment will pose a great challenge to the dynamic resource. So it is mutual exclusion with the limited resources of the aerial MEC system. Therefore, the resource cooperation approach in the aerial MEC network to execute task assignment based on a coalition formation rule needs to be further investigated. To clearly illustrate the difference between the resource coalition cooperation approach and the related works in this field, a comparison in terms of aerial MEC network architecture and optimization objective are provided in Table I. If the performance metric satisfies the corresponding attribute, it is marked with $\checkmark$.
@@
With respect to the system model, most of the aforementioned use one or more UAVs to provide resources in aerial MEC networks. However, a UAV’s resources are limited, making it challenging to support the server requirements of large-scale tasks from MEDs. Although ground base station can provide auxiliary resources, they are often unavailable in emergency scenarios such as military attacks or natural disasters. To address this challenge, some researchers concentrate on the integration of high-altitude platform (HAP) into aerial MEC systems [17, 18]. By establishing effective cooperation between UAVs and HAP, less additional U2U interaction is required when generating resource cooperation strategies. Inspired by the motivation, the emerging technique and practical scenarios, this paper considers DT-empowered task assignment with resource coalition cooperation in an aerial MEC network consisting of multiple users, UAVs equipped with edge servers, and an airship equipped with a more powerful cloud server, i.e., HAP. In this case, our goal is to reduce energy consumption by jointly optimizing the task assignment, energy consumption, communication bandwidth, and the resource utilization. For the sake of clarity, the main contributions of this work are listed as follows
@@
The rest of the section-wise paper outline is detailed as follows. Section II presents a brief synopsis of recent literature in this area. Section III-A introduces the system model. The question statement of the TU coalition is presented in Section IV. A novel TU coalition game algorithm is given in Section V. Section VI analyzes the performance of the coalition game. The simulation results are discussed in Section VII and conclude this paper in Section VIII.
@@ II-C Application of GM in Edge Intelligence Network
To the best of our knowledge, this is the first time we have researched DT-empowered task assignment by coalition game with a GM in an aerial MEC network. Specifically, the system model can monitor the status of the network in real-time and use the method of resource coalition cooperation to optimize task assignment and execution efficiency.
III System Architecture and Model
In this section, we first propose the aerial MEC network architecture for task assignment. Then, we formulate the DT model, communication model and energy consumption model.
III-A System Architecture
This system is designed to use high-altitude airships with high computing power, large caches, and strong communications to assist multi-UAV in completing resource optimization to adapt to MED’s task assignment requests. We build DT models of UAVs in an airship to optimize the resource cooperation of aerial MEC network. With the help of DT, the system model designed in this paper is shown in Fig. 2, which is composed of an application plane, a physical plane, and a virtual plane. The model parameters are shown in Table II. Next, we describe the system model as follows.
@@ III-A System Architecture
III-B DT Model
In this paper, the communication between MED and UAV (M2U) is designed to obtain the MED task parameters to build a DT model. In addition, the communication between UAV and airship (U2A) is designed to transmit the DT modeling parameters of MED and UAV. The above communication process all adopts the communication method of time division multiple access. To simplify the analysis, this paper defines the system model as a synchronous periodic time slot system, and the duration of each time slot $t$ is $\tau$. Let the MED set be $M=\left\{{1,2,\cdots,i,\cdots,m}\right\}$. $m$ means the maximum sequence number of MED. The $i$th MED is represented as $MED_{i}$. Then the DT of $i$th MED in time slot $t$ is
$DT^{t}_{MED_{i}}=\{l^{t}_{MED_{i}},s^{t}_{i},\vartheta^{t}_{i},p^{t}_{i,tr},\tau^{t}_{i}\},$ (1)
where $l^{t}_{MED_{i}}$ represents coordinate $(x^{t}_{i},y^{t}_{i},z^{t}_{i})$ of $MED_{i}$. $s^{t}_{i}$ is the amount of tasks that $MED_{i}$ requests to assign within the time slot $t$. $\vartheta^{t}_{i}$ represents the computing resource required to process the amount of tasks per unit, that is, computational complexity. $p^{t}_{i,tr}$ represents the transmission power of $MED_{i}$ for task assignment. $\tau^{t}_{i}$ means the time slot constraint of task, i.e., the maximum time allowed to complete the task of $MED_{i}$, which is $\tau^{t}_{i}\leq\tau$.
The set of UAV is denoted as $V=\left\{{1,2,\cdots,j,\cdots,n}\right\}$. $n$ means the maximum sequence number of UAV. Let $UAV_{i}$ be the $i$th UAV, DT of $UAV_{j}$ within time slot $t$ is given as follows
$DT^{t}_{UAV_{j}}=\{l^{t}_{UAV_{j}},b^{t}_{j},f^{t}_{j},c^{t}_{j},p^{t}_{j,h}\},$ (2)
@@ III-B DT Model
TABLE II: List of parameters
@@
$\mu$ The utilization of computing resource
III-C Communication Model
To illustrate the communication capacity between a MED and an UAV coalition (i.e., One-to-many assigning, see Section IV-A), we assume that one $MED_{i}$ interacts with an UAV coalition in each time slot $t$. The coordinated multiple points transmission technology based on orthogonal frequency division multiple access is introduced in the UAV coalition. And allows UAV to reuse spectrum resources with a total bandwidth of $B^{t}_{max}$ to reduce communication interference. At the same time, other UAVs and MED outside the UAV coalition are in energy-saving mode. Further, the communication capacity between $MED_{i}$ and $UAV_{j}$ in the time slot $t$ is given as follows
$C^{t}_{ij}=b^{t}_{j}\log\left(1+\dfrac{|g^{t}_{ij}|p^{t}_{i,tr}}{\sigma^{2}}\right),$ (3)
@@ III-C Communication Model
III-D Energy Consumption Model
@@ IV The Question Statement of Task Assignment
In this section, we first introduce the task assignment problem in the aerial MEC network and discuss its differences compared with conventional task offloading in MEC networks. Then, the resource coalition cooperation algorithm is proposed.
IV-A Definition of Task Assignment
The cooperation of resources from various devices is a fundamental problem in the MEC network. The basic idea of task assignment is to offload the tasks to the appropriate computing devices, which is to improve the utilization of resources and the efficiency of task execution. However, in order to find the optimized task assignment strategy, we are mapping the multiple UAVs to a resource pool. Compared with the existing task offloading in MEC systems, the task assignment in this paper has its own unique characteristics in the aerial MEC network.
Peer-to-peer pooling: All the nodes equipped with various resources in the aerial MEC network are regarded as equal, which builds a resource pool among the nodes. Furthermore, the resources of the nodes can be used through cooperative games, and their resource capabilities will be uniformly scheduled.
Low-to-high scheduling: The nodes in the aerial MEC network are composed of end users (i.e., MEDs), edge servers (i.e., UAVs), and cloud servers (i.e., airship). The tasks in our proposed network can be executed in low-to-high manner according to task request: 1) one-to-one executing; 2) one-to-many assigning. In this paper, we focus on scheduling resources on multiple nodes to complete the tasks.
One-to-many assigning: In existing task offloading strategies, the tasks are usually offloaded to a powerful edge server or a remote cloud server for increasing computing efficiency (i.e., one-to-one executing). However, in the proposed aerial MEC network, the tasks can be assigned to many edge servers.
@@ IV-C TU Coalition Utility
$\begin{array}[]{ll}F1:&\underset{s^{t}_{j},b^{t}_{j}}{\text{max}}~{}~{}~{}U^{t}_{g_{k}}\\ \text{ s.t.: }&C_{1}:0\leq b^{t}_{j}\leq b^{t}_{j,max},\\ &C_{2}:\Sigma_{j\in g_{k}}b^{t}_{j}\leq B^{t}_{max},\\ &C_{3}:0\leq s^{t}_{j}\leq c^{t}_{j,max},\\ &C_{4}:\Sigma_{j\in g_{k}}s^{t}_{j}\leq s^{t}_{i},\\ &C_{5}:T^{t}_{gk}\leq\tau^{t}_{i},\end{array}$ (9)
@@ IV-D Coalition Participant Utility
$u_{j}^{t}=U^{t}_{g_{k}}\frac{s_{j}^{t}}{\sum_{j\in g_{k}}s_{j}^{t}}-\alpha E_{j,cp}^{t}-\beta E_{j,h}^{t},$ (10)
@@
$\begin{array}[]{ll}F2:&\underset{f^{t}_{j}}{\text{max}}~{}~{}~{}u^{t}_{j}\\ \text{ s.t.: }&\mathrm{C}_{6}:0\leq f_{j}^{t}\leq f^{t}_{j,max},\\ &\mathrm{C}_{7}:T_{j,cp}^{t}\leq\tau_{i}^{t}-T_{j,tr}^{t},\end{array}$ (11)
@@
In our proposed method, the UAV acts as an aerial base station for providing computing services. Therefore, the utilization of the computing resources of the UAV must be considered when obtaining the maximum participant utility. The utilization of the computing resource $mu$ by each participant in the coalition is designed as follows
$\mu=\frac{\tilde{f}_{j}^{t}}{f^{t}_{j,max}},$ (12)
where $\tilde{f}_{j}^{t}$ is the actual value of the computing resource of the UAV.
@@ IV-E Problem Conversion and Merging
$\mathrm{C}_{8}:s_{j}^{t}\leq\tau_{i,\max}^{t}/\left(\frac{1}{C_{ij}^{t}}+\frac{\vartheta_{j}^{t}}{f_{j,\text{ max }}^{t}}\right),$ (14)
@@
$\begin{array}[]{ll}F3:&\underset{s_{j}^{t},b_{i}^{t}}{\text{max}}~{}~{}~{}U^{t}_{g_{k}}\\ \text{ s.t.: }&\mathrm{C}_{1},\mathrm{C}_{2},\mathrm{C}_{3},\mathrm{C}_{4},\mathrm{C}_{8},\mathrm{C}_{9}.\end{array}$ (16)